The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31 Amount Total sales revenue Selling price per pair of skis Variable selling expense per pair of skis Variable administrative expense per pair of skis Total fixed selling expense Total fioxed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases $1,419,000 $430 49 17 S 160,000 S 105,000 S 80,000 S 100,000 s 290,000 Required: 1. Prepare a traditional income statement for the quarter ended March 31 The Alpine House, Inc. Traditional Income Statement 1,419,000 Sales Cost of goods sold Gross margin Selling and administrative expenses Selling expenses Administrative expenses Net operating income
media%2F945%2F9455d3c3-6910-4024-aefc-36
Answer #1
Units sold 3300 =1419000/430
1
Traditional income statement
Sales 1419000
Cost of goods sold 270000 =290000+80000-100000
Gross margin 1149000
Selling and administrative expenses:
Selling expenses 321700 =160000+(3300*49)
Administrative expenses 161100 =105000+(3300*17)
482800
Net operating income 666200
2
Contribution format income statement
Sales 1419000
Variable expenses:
Cost of goods sold 270000
Selling expenses 161700 =3300*49
Administrative expenses 56100 =3300*17
487800
Contribution margin 931200
Fixed expenses:
Selling expenses 160000
Administrative expenses 105000
265000
Net operating income 666200
3
Contribution of each pair of skis 282 =931200/3300